OpenStudy (anonymous):

please help me!! how can 1/5x-2=1/3x+8 be set up as a system of equations

OpenStudy (jdoe0001):

do you know what a system of equations is?

OpenStudy (anonymous):

yes its two things like y = 3x+2 and y =2x+1

OpenStudy (anonymous):

but I don't understand how to do that one

OpenStudy (jdoe0001):

right thus \(\large \bf \cfrac{1}{5}x-2=\cfrac{1}{3}x+8\implies \begin{cases} y=\cfrac{1}{5}x-2 \\ \quad \\ \bf y=\cfrac{1}{3}x+8 \end{cases}\)

OpenStudy (anonymous):

see thats what I get but my choices arent like that 5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (jdoe0001):

well, if you multiply ... one sec

OpenStudy (anonymous):

multiply what

OpenStudy (jdoe0001):

well... I got rid of the rationals, doesn't give me your choices though

OpenStudy (jdoe0001):

I gather what I would do is, find "x" and "y" and check it against the given choices

OpenStudy (jdoe0001):

I think it'd be the simpler way to go about it

OpenStudy (anonymous):

Okay